# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v2

@@ 3 A Case for PiM-based Timing Attacks
The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any of their vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work investigates and evaluates the security challenges of new and emerging PiM architectures. After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions in the forms of processing-near-memory (PnM) and processing-using-memory (PuM) exposes the internal states of the main memory (e.g., row buffer states of DRAM banks) by providing direct access to the user space applications. Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [66, 67, 68], being either inapplicable or performance-inefficient in PiM architectures.
@@ 3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already cached in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands, $PRE$ to precharge the open row and $ACT$ to fetch the target row to the row buffer. Therefore, it is possible for a user space application to infer whether a row is accessed recently by measuring the time it takes to access the row. To demonstrate the latency gap between different states of the row buffer, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts, and we measure the latencies observed by the user space application. Figure 2 shows the breakdown of memory access latencies categorized into two classes: row buffer hits and conflicts.
Figure 2: The distribution of memory access latencies.
We observe that the average latency gap between a row buffer conflict and a row buffer hit is 74 CPU cycles. Since this latency gap is large enough to detect, it is possible for user space applications that share the main memory to exploit the row buffer as a covert channel to exchange messages. Our results align with the observations of prior works [70].
@@ 3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on applications that share main memory and establish covert channels to exchange messages between different processes. However, these covert-channel attacks are limited in throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests that are eventually sent to main memory.
There are several ways to bypass the cache hierarchy in traditional systems by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by direct memory access (DMA) and remote direct memory access (R-DMA) engines, and (iv) non-temporal memory hints (e.g., movnti). However, when used in a main memory-based timing covert channel attack, these methods do not provide fast and reliable communication.
@@
We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties.
First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the latency of writing back a cache line to the main memory on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications.
Second, eviction sets require a large number of load requests proportional to the number of ways in the caches, which scales with modern designs. Therefore, the attacks that use them have limited throughput. While the ISA guarantees the instructions used to create the eviction sets, the eviction set is not guaranteed to evict the target cache line (due to cache characteristics such as cache replacement policy, cache prefetchers etc.).
Third, DMA and R-DMA engines incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example attack using this primitive[103] showcases a covert channel attack based on memory contention. However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput these covert channels provide is limited (in the order of Kb/s).
Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach. Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches.
In contrast, PiM architectures provide fast and reliable direct memory access to user space applications that can be used to establish a covert-channel attack.
@@ 3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the latencies induced by deep cache hierarchies (i.e., cache access and cache eviction latencies) on communication throughput, we simulate a realistic system (described in Section 5.1) that establishes a state-of-the-art covert channel [70].
@@
Attack Scenarios. We simulate two attacks. First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets. To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line and then, it fetches the target row into the row buffer using an additional memory request. The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache. Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads. Second, we evaluate a direct memory access attack that can directly access the main memory. The direct memory access attack issues a memory request to fetch the data to the row buffer and does not need any cache access or eviction steps.
@@
For both attacks, we measure the time it takes to send a bit through the row buffer including cache access and cache eviction latencies, and with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time needed to propagate bits from the sender to receiver. We analyze the impact of (i) the LLC size, and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
@@
Impact of LLC Size on Communication Throughput. Figure 3 shows the throughput of baseline and direct memory access attacks. The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [104]. To better understand the throughput results, Figure 4 shows the latency of a cache line eviction of caches with corresponding sizes.
Figure 3: Impact of the LLC size on the data leakage throughput of covert channels.
Figure 4: Latency of a cache line eviction in CPU cycles with increasing LLC size.
@@
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27Mb/s$ across all LLC sizes, which is significantly greater than the throughput of up to $2.29Mb/s$ achieved by the baseline attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes due to increasingly expensive cache access and eviction latencies. We observe that with increasing LLC size, the eviction latency grows significantly due to the increase in access latency. We conclude that with increasing LLC sizes, the cache access and eviction latencies grow significantly and significantly hinder the communication throughput of the baseline attack compared to the direct memory access attack which is independent of the LLC size.
Impact of LLC Associativity on Communication Throughput. Figure 5 shows the communication throughput of baseline and direct memory access attacks. The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core). Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache. To better understand this effect, Figure 6 shows the latency of a cache line eviction across increasing numbers of LLC ways. We observe that the baseline attack’s communication throughput decreases significantly with increasing number of LLC ways due to the increased cache eviction latency.
@@
Figure 5: Impact of the number of LLC ways on the data leakage throughput of covert channels.
Figure 6: Latency of evicting a set in CPU cycles with increasing number of LLC ways.
@@
We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache lookup and eviction latencies, which worsen with increasing size and associativity. However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
@@
Adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications. Therefore, it is crucial to carefully consider the security implications of PiM architectures and to develop robust security measures to protect against potential attacks. Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures. We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
4 IMPACT: A High-Throughput Main Memory Timing Attack
@@
This section introduces IMPACT, a set of high-throughput main memory timing attacks that takes advantage of the direct memory access enabled by the emerging processing-in-memory paradigm. IMPACT (i) operates independently of specialized instructions that flush data from the cache hierarchy and (ii) avoids the resource-intensive cache eviction steps typically required for main memory and cache-based timing attacks, achieving higher data leakage throughput than previous memory-based attacks. In this work, we present (1) two variants of IMPACT covert channel attack and (2) an IMPACT side-channel attack on DNA read mapping applications enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions. We evaluate the communication throughput of IMPACT covert channels in Section 5.3 and the leakage throughput of IMPACT side channel in Section 5.4.
4.1 IMPACT Covert-Channel Attacks
4.1.1 PnM-based Covert-Channel Attack
@@
Baseline PnM Architecture. In IMPACT-PnM, we consider PiM-Enabled Instructions (PEI)  [48] as the baseline PnM architecture. PEI expands the instruction set architecture (ISA) of the host processor with PIM-enabled instructions (PEIs), while keeping the existing programming model intact. A PEI is an instruction that can be executed either on the host processor or on the in-memory computation logic which is integrated near each DRAM bank of an HBM stack. PEI can also easily be adopted in other memory technologies since its design does not rely on specific features of the memory technology. We chose PEI as the baseline PnM architecture as it introduces a simple and effective model (i.e., offloading computation at the instruction granularity) for PnM, and our attack vector can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
@@
Figure 7 depicts the organization of PEI. PEI architecture consists of two key components. First, PEI integrates a PEI Computation Unit (PCU) near each DRAM bank  1 and inside the CPU  2. In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, PEI adds a PEI Management Unit (PMU)  3 integrated with the last-level cache of the host system to (i) monitor the data locality of application regions and (ii) map each region to host-side or memory-side PCUs based on their data locality. This way a PEI can be selectively executed on the host processor PCU (instead of the in-memory PCU) when large on-chip caches are beneficial for its execution. As mentioned in [105], if a locality monitor entry is allocated by a PIM operation, it is often too aggressive to consider it as having high locality on its first hit in the locality monitor. Therefore, they propose adding an ignore flag per entry to ignore the first hit to an entry allocated by a PIM operation. Such a design allows us to easily bypass the locality monitor even with a small virtual address range in our disposal.
Figure 7: Overview of PEI’s architecture.
Attack Overview. The PnM-based covert channel (IMPACT-PnM) leverages the direct memory access provided by PiM-enabled instructions, which bypasses the cache hierarchy. IMPACT-PnM exploits the timing difference between the latency of a DRAM row buffer hit and a conflict to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference (i.e., row buffer hits or conflicts) in the row buffer of a DRAM bank. We assume the sender and the receiver co-locate data in the same DRAM banks. [^1] If PiM is the main memory, then the sender can co-locate its data with the receiver process’s data by using memory massaging techniques, as used in many prior works [90, 70, 106, 107]. Two processes can share data itself in various cases, such as databases and websites running on browsers. In this attack, we encode logic-1 as interference in the DRAM bank, and logic-0 as no interference.
@@
Figure 8 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
Step 1. The receiver issues PEI operations to initialize each DRAM bank by activating a predetermined row (Line 18). It is necessary that the cache lines of the activated row are not used before in order to bypass the locality monitor (PMU) that is employed in PEI-based systems. Then, the sender and the receiver synchronize on a barrier (Lines 5&20) which ensures that they start the communication turn at the same time.
Step 2. The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted. To transmit a 0, as shown in Fig. 8 ( 1) the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains activated in the row buffer  2). To transmit a 1, as shown in Fig. 8 ( 3) the sender activates a row that is different from the one activated by the receiver to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4). At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed before the receiver starts probing the DRAM banks and then signals the receiver to start probing the DRAM banks (Line 15).
Step 3. The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25). Then, the receiver measures the time it takes to complete the request (Line 27). If the latency of executing the PEI exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6). Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8). The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.

### v3

@@ 3 A Case for PiM-based Timing Attacks
The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions provides user space applications with fast and reliable direct access to main memory. Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [64, 65, 66], being either inapplicable or performance-inefficient in PiM architectures.
@@ 3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands to precharge the open row and fetch the target row to the row buffer. Therefore, a user space application can infer whether a row has been accessed recently by measuring the time it takes to access the row. To demonstrate the latency gap between different row buffer states, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts.
We observe that a row buffer conflict takes 74 CPU cycles more than a hit, which is large enough to detect. So, a user space application can exploit the row buffer as a timing covert and side channel. Our results align with the observations of prior works [68, 77, 78].
@@ 3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory. However, these covert-channel attacks have limited throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests.
There are several ways to bypass the cache hierarchy by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by (remote) direct memory access, (R)DMA, engines, and (iv) non-temporal memory hints (e.g., movnti). However, no attack primitive provides both fast and reliable communication.
@@
We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties. First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications. Second, eviction sets require many load requests proportional to the number of ways in the caches, which scales with modern designs. Therefore, the attacks that use them have limited throughput. While the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other design choices that are not transparent to the attacker. Third, DMA and R-DMA engines incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example attack using this primitive[91] showcases a covert-channel attack based on memory contention. However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s). Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach. Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches. In contrast, PiM operations provide fast and reliable direct memory access to user space applications, which can be exploited for high-throughput covert- and side-channel attacks.
@@ 3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a realistic system (described in §5.1), establishing a state-of-the-art covert channel [68].
@@
Attack Scenarios. We simulate two attacks. First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets. To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line, and then it fetches the target row into the row buffer using an additional memory request. The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache. Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads. Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
@@
For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time needed to propagate bits from the sender to the receiver. We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
@@
Impact of LLC Size on Communication Throughput. Fig. 2 shows the impact of the LLC size on (i) the throughput of baseline and direct memory access attacks and (ii) the latency of a cache line eviction. The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [92].
@@
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive. With increasing LLC size, the cache access latency increases and, thus, amplifies the cost of eviction, which requires multiple cache accesses.
@@
Figure 2: Impact of the LLC size on the data leakage throughput of covert channels and the eviction latency.
Impact of LLC Associativity on Communication Throughput. Fig. 3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks and (ii) the eviction latency of a cache line. The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core). Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache. We observe that the baseline attack’s communication throughput decreases significantly with an increasing number of LLC ways due to the higher cache eviction latency.
Figure 3: Impact of the number of LLC ways on the data leakage throughput of covert channels and the eviction latency.
@@
We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity. However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
@@
The adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications. Therefore, it is crucial to carefully consider the security implications of PiM architectures and develop robust security measures to protect against potential attacks.
4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
@@
This section introduces IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by the emerging processing-in-memory paradigm. IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. In this work, we present two covert-channel and one side-channel attacks enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
4.1 PnM-based IMPACT Covert-Channel Attack
@@
Baseline PnM Architecture. PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture of the host processor with PIM-enabled instructions (PEIs). PEIs can be executed either on the host processor or on the in-memory computation logic. We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
@@
The baseline PEI architecture has two key components. First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU. In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, the PEI architecture has a PEI Management Unit (PMU) that monitors data locality in application regions and decides where to map each region (either to host-side or memory-side PCUs). This way, PEI architecture can execute regions with high locality on the host side to benefit from caches. PEI adds an ignore flag per locality monitor entry to ignore the first hit in the locality monitor [93]. This is because if a locality monitor entry is allocated for a PIM operation, it is often too aggressive to consider the operation as having a high locality on its first hit. This bit allows IMPACT-PnM to easily bypass the locality monitor even with a small virtual address range at its disposal.
Attack Overview. Before the attack, the sender and the receiver co-locate their data in the same DRAM banks. To achieve this, one process uses memory massaging techniques to place its data in the same bank as the other process in main memory, as used in many prior works [76, 68, 94, 95].
IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank. In this attack, we encode logic-1 as interference in the DRAM bank (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
@@
Fig. 4 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
Figure 4: Flow of the PnM-based Covert Channel Attack.